# GUI-Bee [Uncaptioned image]     : Align GUI Action Grounding to Novel Environments via Autonomous Exploration

### v1

@@ 1 Introduction
GUI action grounding maps natural language instructions to specific executable elements or locations on a GUI screen. It is critical as being a key step adopted by GUI agents to assist humans in complex digital environments, where GUI agents usually generate step-by-step instructions and then rely on the action grounding to locate corresponding executable GUI elements(Agashe et al., 2024; Zheng et al., 2024a; Koh et al., 2024a). As the planning ability of MLLMs shines, their relatively weak zero-shot grounding performance becomes a key bottleneck in building strong GUI agents (Zheng et al., 2024b), and GUI action grounding has become one of the focus of recent specialized model development efforts (Gou et al., 2024; Cheng et al., 2024; Liu et al., 2024; Chen et al., 2024).
@@ 2.1 GUI Grounding with MLLMs
As GUI agents evolve from text-based (Sodhi et al., 2024; Zhou et al., 2023) to multimodal systems (Deng et al., 2023; Koh et al., 2024a; Xie et al., 2024), GUI action grounding—linking natural language queries to GUI elements—has become a key challenge and shifted from being part of GUI agent tasks to an independent focus. Starting from GUI action grounding solutions (Zheng et al., 2024b; Koh et al., 2024b) that zero-shot MLLMs with the SoM methods (Yang et al., 2023), recent works on GUI grounding models have put efforts into emphasizing generalization. Pioneering works like SeeClick (Cheng et al., 2024) introduced visual-only grounding that is easier generalized to different platforms, avoiding the limitations of platform-specific structured text, while studies such as GUICourse (Chen et al., 2024) expanded beyond grounding executable GUI elements to general GUI element. More recent works further address challenges introduced by screenshots like high resolution and interleaved text and images by enhancing model designs (Gou et al., 2024; Lin et al., 2024; You et al., 2024) and leveraging large-scale in-domain training data (Wu et al., 2024; Liu et al., 2024). However, these previous works overlook that GUI action grounding queries are often environment-dependent. In our work, we boost GUI action grounding models by continuously fine-tuning them with environment-specific data collected by our GUI-Bee agent. Our method can be applied on top of all other GUI grounding models and significantly improve their performance.
@@ 2.2 In-Context Learning
In-context learning (ICL) refers to the method of adapting models to new tasks by providing context (Brown, 2020; Chan et al., 2022; Wang et al., 2023). By including examples directly in prompts, ICL allows large language models (LLMs) to generalize to unseen tasks (Garg et al., 2022; Pan, 2023; Wei et al., 2023). Prior works have explored applying ICL to reinforcement learning (RL) either with model training involved (Laskin et al., 2022; Lee et al., 2024; Xu et al., 2022) or by directly leveraging pre-trained LLMs (Krishnamurthy et al., 2024; Monea et al., 2024). We propose the Q-value-incentive In-context Reinforcement Learning (Q-ICRL), which also utilizes pre-trained LLMs but distinguishes itself by using ICL to predict state-action values. Our approach combines the adaptability of LLMs with RL’s optimization-driven structure, enabling efficient action selection in the GUI environment exploration.
@@ 3.1.2 Challenges in the Exploration
The exploration process faces several challenges, starting with identifying valid actions within a noisy action space. The set of action candidates $A_{\text{env}}(i^{t})$, obtained from the environment, is defined as $a^{t}\in A_{\text{env}}(i^{t})=\{a^{t}_{1},a^{t}_{2},\dots,a^{t}_{n}\},$ and can originate from sources such as a Document Object Model (DOM) tree or a separate module $M_{\text{action}}(i^{t})$, like OmniParser (Lu et al., 2024). However, $A_{\text{env}}(i^{t})$ often includes invalid actions targeting non-executable elements, requiring the agent to discern optimal actions that are both valid and meaningful for exploration.
@@ 3.1.3 Q-value-Incentive In-Context Reinforcement Learning (Q-ICRL)
We propose the Q-value-Incentive In-Context Reinforcement Learning (Q-ICRL) method, outlined in Algorithm 1, to maximize the accumulated rewards in the exploration. Q-ICRL quantifies the future rewards of executing actions $a^{t}$ at $S^{t}$ with a Q-value function $Q(S^{t},a^{t})$. Unlike traditional Q-learning, Q-ICRL is a training-free algorithm as the $Q$ is mainly memory-based, which is detailed in the following subsection. To select an action $a^{t}$ at state $S^{t}$, Q-values $Q(S^{t},a^{t}_{i})$ are first used as weights to sample a subset $A^{\prime}_{\text{env}}(i^{t})$ of length $H$ from the action space $A_{\text{env}}(i^{t})$. Then, an MLLM with in-context learning is employed to identify the most promising action $a^{t}$ from $A^{\prime}_{\text{env}}(i^{t})$, guided by the Q-value function. This process involves two steps: ExampleActionIdentification and InContextScoring, both detailed in the next subsection. This method ensures a balance between exploration and exploitation. Once an action $a^{t}$ is selected, it is executed in the environment, leading to an updated exploration graph $G^{t+1}=G^{t}\oplus(a^{t},i^{t+1})$, where $\oplus$ denotes adding a new node and edge to the graph. The Q-value function is updated accordingly, as described in 3.1.1.
@@
Given $a^{t}_{i}$ in $A_{\text{env}}(i^{t})$, the $Q(S^{t},a^{t}_{i})$ should be relatively low if executing it leads to a $i^{t+1}$ that is repeated in $S^{t}$, however, as mentioned in 3.1.2, the $i^{t+1}$ unknown at $S^{t}$. To overcome this problem, We propose a Q-value function that approximates $Q(S^{t},a^{t}_{i})$ with $Q(S^{x},a^{t}_{i})$ when there is a $S^{x}$, $x\in[1,\dots,t-1]$ with the same $a^{t}_{i}$ was executed. As for $a^{t}_{i}$ that is never executed in the past states, $Q(S^{t},a^{t}_{i})$ is set at a default value of 100.
@@
To offset the potential error in the Q-value function, we leverage the knowledge in the MLLM to determine the most suitable action. After $A^{\prime}_{\text{env}}(i^{t})$ is sampled, we let the MLLM to predict $\hat{Q}(S^{t},a^{t}_{x})$ for each sampled candidate action $a^{t}_{x}\in A^{\prime}_{\text{env}}(i^{t})$. $\hat{Q}(a^{t}_{x})$ is an MLLM-approximation of the $Q(S^{t},a^{t}_{x})$, and the in-context learning is employed. Specifically, the MLLM is provided with the current state $S^{t}$ and two example actions $(a_{\text{eg1}},a_{\text{eg2}})$ along with their corresponding Q-values $(Q(S^{t},a_{\text{eg1}}),Q(S^{t},a_{\text{eg2}}))$. The example actions are chosen from $A_{\text{env}}(i^{t})$ with $(Q(S^{t},a_{\text{eg1}}),Q(S^{t},a_{\text{eg2}}))\neq 100$ while having the corresponding element $e_{\text{eg1}}$ and $e_{\text{eg2}}$ most similar to the $e^{t}_{x}$. We develop a GUI Element Fuzzy Visual Matching module, detailed in Appendix A, to identify the visual similarity between element visuals. Next, we mark $(a_{\text{eg1}},a_{\text{eg2}})\text{ and }a^{t}_{x}$ on the screenshot of $i^{t}$ as the visual input to the MLLM with the Set-of-Mark method (Yang et al., 2023). Figure 3 shows an example of this process, and the full prompts are detailed in Appendix D. If no suitable $a_{\text{eg}}$ is found, the example-related content is omitted from the prompt. Finally, the action $a^{t}_{x}$ with the highest predicted $\hat{Q}(a^{t}_{x})$ is selected as the action $a^{t}$ to be executed.
@@ 3.2 Autonomous Data Annotation with GUI-Bee
After the exploration, for each edge $a^{t}$ in the $G$, the connected nodes $(i^{t},i^{t+1})$ is sent to the MLLM to generate $u^{t}$, a list of queries serving as action grounding queries for the target element $e^{t}$ in ${i}^{t}$. We carefully design the prompt, detailed in Appendix B, to guide the MLLM in generating these queries. The queries involve both "what is currently visible" on the screen and "what will appear" after interacting with GUI elements, which we refer as System 1 and System 2 grounding queries, inspired by Kahneman (2011). Building on the multi-lens prompting method (Fan et al., 2024), we create separate visual prompts, or lenses: two capture the full screens of $i^{t}$ and $i^{t+1}$ while one isolates $e^{t}$ to ensure high-quality outputs. The center of $e^{t}$ is then sampled as the target point $p^{t}$, and the resulting data ($u^{t},{i}^{t},p^{t}$) is used to fine-tune GUI action grounding models.
@@ 4 NovelScreenSpot Benchmark
NovelScreenSpot simulates real-world GUI model deployment scenarios, where App owners want to evaluate GUI action models on their own specific GUI environments, potentially novel environments that the models are not previously trained on. Therefore, unlike existing benchmarks that emphasize diversity across many environments, such as the ScreenSpot (Cheng et al., 2024), NovelScreenSpot instead provides a greater data variation within each environment. Notably, the NovelScreenSpot includes a large number of queries, around one-third of the total benchmark, focusing on interaction outcomes. These queries require environment-specific knowledge and hardly exist in the existing benchmarks. The five GUI environments in the NovelScreenSpot are three offline websites from the VisualWebArena (Koh et al., 2024a)—Shopping, Classifieds (a second-hand marketplace), and Reddit (an online forum)—and two online websites, Photoshop-web and Eventbrite. These environments vary greatly in style, with Photoshop-web dominated by professional icons, Shopping and Classifieds emphasizing images, and Reddit and Eventbrite focusing more on textual content.
@@ 5.1 Alignment to Novel GUI Environments
We employ the GUI-Bee agent to explore the five environments in NovelScreenSpot, conducting up to 400 exploration steps per environment with three candidate actions sampled at each step. To ensure diverse screen data, the exploration is repeated three times per environment at varying screen resolutions. In this work, we adopt GPT-4o (OpenAI, 2024) as the multimodal large language model (MLLM), though other MLLMs can also be integrated into the framework. The resulting exploration statistics are summarized in Table 1. Using the data generated from these explorations, we continuously fine-tune four GUI grounding models: SeeClick (Cheng et al., 2024), Qwen-GUI (Chen et al., 2024), and UIX-7B (Liu et al., 2024) to adapt to novel environments within NovelScreenSpot. Additionally, for the alignment of Qwen-GUI and UIX-7B, we leverage both vision-only and Vision+A11y data, as these models were previously fine-tuned to accommodate similar formats. Further details on exploration configurations, data formatting, and fine-tuning settings are provided in Appendix E.
@@
We present the model performance before and after alignment to novel GUI environments in Table 2. The models are either continuously fine-tuned using exploration data collected by the GUI-Bee agent or, as a baseline, training data from Multimodal-Mind2Web (Zheng et al., 2024a). Our results show that data collected by GUI-Bee significantly outperforms the baseline in improving model performance across all tested models. We attribute this to the environment-specific knowledge captured by the GUI-Bee agent, which contrasts with Multimodal-Mind2Web’s design for generalization across environments. Notably, the UIX-7B model achieves the most substantial accuracy gains and the highest overall accuracy after alignment. Furthermore, models fine-tuned in the vision+A11y GUI action grounding setting demonstrate greater improvements. This is likely because the original GUI grounding models were not trained with accessibility (A11y) tree information, and the inclusion of A11y data provides additional screen context, enabling more accurate predictions. These findings underscore the value of the diverse formats in GUI-Bee’s collected data, which incorporate contextual information such as accessibility data to enhance model performance.
@@
We extend our evaluation by following Gou et al. (2024) to leverage the test data from Multimodal-Mind2Web (Zheng et al., 2024a) for assessing the alignment of GUI grounding models to novel environments. This test data simulates GUI grounding tasks guided by MLLM-planned instructions in GUI agent applications. Multimodal-Mind2Web data is derived from real action sequences across diverse GUI environments, corresponding to high-level, real-world GUI tasks. The grounding queries consist of element descriptions generated by MLLMs, without explicitly referencing coordinates. As the Eventbrite environment is the only one from Multimodal-Mind2Web included in NovelScreenSpot, we adopt corresponding data from the Multimodal-Mind2Web test set to evaluate models that we have aligned to the Eventbrite environment. Results reveal that models continuously fine-tuned with GUI-Bee data achieve significantly greater performance improvements compared to baseline models using Multimodal-Mind2Web training set data. This further validates the effectiveness of our proposed method.
@@ 5.2 Exploration Efficiency Evaluation
To further evaluate the efficiency of our GUI-Bee agent in exploring and generating data within GUI environments, we task it, along with two baseline agents, to explore three offline GUI environments: Shopping, Classifieds, and Reddit—from the Visual Web Arena (Koh et al., 2024a). These environments are reset to identical initial states at the beginning of each exploration, ensuring that all agents start from the same conditions and face equivalent challenges.
@@
To compare with our GUI-Bee agent, we introduce two baseline agents: the In-Context Reinforcement Learning (ICRL) method and the random exploration strategy. The ICRL agent is an ablated version of our GUI-Bee agent. It uses the same RL state representation as the GUI-Bee agent but does not build or utilize the Q-value function. Instead, it directly leverages the MLLM with in-context learning to select the next action on the GUI screen. The in-context example is randomly chosen from the actions recorded in the exploration graph. The random agent follows a purely stochastic strategy, selecting actions randomly from the set of candidate actions on each screen. All agents are constrained to the same maximum number of action steps to enable a fair comparison.

### v2

@@ Abstract
GUI-Bee [Uncaptioned image]     : Align GUI Action Grounding to Novel Environments via Autonomous Exploration
@@
22footnotetext: Co-advising. This work was partly performed when the first author interned at Adobe Research.
@@ 1 Introduction
GUI action grounding maps natural language instructions to specific executable elements or locations on a GUI screen. It is critical as being a key step adopted by GUI agents to assist humans in complex digital environments, where GUI agents usually generate step-by-step instructions and then rely on the action grounding to locate corresponding executable GUI elementsAgashe et al. (2024); Zheng et al. (2024b); Koh et al. (2024a). While MLLMs excel in planning, their weak zero-shot grounding performance limits the effectiveness of GUI agents Zheng et al. (2024a). Consequently, GUI action grounding has become a focus of recent specialized model development efforts Gou et al. (2024); Cheng et al. (2024); Liu et al. (2024); Chen et al. (2024).
@@ 2.1 GUI Grounding with MLLMs
As GUI agents evolve from text-based Sodhi et al. (2024); Zhou et al. (2023) to multimodal systems Deng et al. (2023); Koh et al. (2024a); Xie et al. (2024), GUI action grounding—linking natural language queries to GUI elements—has become a key challenge and shifted from being part of GUI agent tasks to an independent focus. Starting from GUI action grounding solutions Zheng et al. (2024a); Koh et al. (2024b) that zero-shot MLLMs with the SoM methods Yang et al. (2023), recent works on GUI grounding models have put efforts into emphasizing generalization. Pioneering works like SeeClick Cheng et al. (2024) introduced visual-only grounding that is easier generalized to different platforms, avoiding the limitations of platform-specific structured text, while studies such as GUICourse Chen et al. (2024) expanded beyond grounding executable GUI elements to general GUI element. More recent works further address challenges introduced by screenshots like high resolution and interleaved text and images by enhancing model designs Gou et al. (2024); Lin et al. (2024); You et al. (2024) and leveraging large-scale in-domain training data Wu et al. (2024); Liu et al. (2024). However, these previous works overlook that GUI action grounding queries are often environment-dependent. In our work, we boost GUI action grounding models by continuously fine-tuning them with environment-specific data collected by our GUI-Bee agent. Our method can be applied on top of all other GUI grounding models and significantly improve their performance.
@@ 2.2 In-Context Learning
In-context learning (ICL) refers to the method of adapting models to new tasks by providing context Brown (2020); Chan et al. (2022); Wang et al. (2023). By including examples directly in prompts, ICL allows large language models (LLMs) to generalize to unseen tasks Garg et al. (2022); Pan (2023); Wei et al. (2023). Prior works have explored applying ICL to reinforcement learning (RL) either with model training involved Laskin et al. (2022); Lee et al. (2024); Xu et al. (2022) or by directly leveraging pre-trained LLMs Krishnamurthy et al. (2024); Monea et al. (2024). We propose the Q-value-incentive In-context Reinforcement Learning (Q-ICRL), which also utilizes pre-trained LLMs but distinguishes itself by using ICL to predict state-action values. Our approach combines the adaptability of LLMs with RL’s optimization-driven structure, enabling efficient action selection in the GUI environment exploration.
@@ 3.1.2 Challenges in the Exploration
The exploration process faces several challenges, starting with identifying valid actions within a noisy action space. The set of action candidates $A_{\text{env}}(i^{t})$, obtained from the environment, is defined as $a^{t}\in A_{\text{env}}(i^{t})=\{a^{t}_{1},a^{t}_{2},\dots,a^{t}_{n}\},$ and can originate from sources such as a Document Object Model (DOM) tree or a separate module $M_{\text{action}}(i^{t})$, like OmniParser Lu et al. (2024). However, $A_{\text{env}}(i^{t})$ often includes invalid actions targeting non-executable elements, requiring the agent to discern optimal actions that are both valid and meaningful for exploration.
@@ 3.1.3 Q-value-Incentive In-Context Reinforcement Learning (Q-ICRL)
We propose the Q-value-Incentive In-Context Reinforcement Learning (Q-ICRL) method, outlined in Algorithm 1, to maximize the accumulated rewards in the exploration. Q-ICRL quantifies the future rewards of executing actions $a^{t}$ at $S^{t}$ with a Q-value function $Q(S^{t},a^{t})$. Unlike traditional Q-learning, Q-ICRL is a training-free algorithm as the $Q$ is mainly memory-based, which is detailed in the following subsection. To select an action $a^{t}$ at state $S^{t}$, Q-values $Q(S^{t},a^{t}_{i})$ are first used as weights to sample a subset $A^{\prime}_{\text{env}}(i^{t})$ of length $H$ from the action space $A_{\text{env}}(i^{t})$. Then, an MLLM with in-context learning is employed to identify the most promising action $a^{t}$ from $A^{\prime}_{\text{env}}(i^{t})$, guided by the Q-value function. This process involves two steps: ExampleActionIdentification and InContextScoring, both detailed in the next subsection. This method ensures a balance between exploration and exploitation. Once an action $a^{t}$ is selected, it is executed in the environment, leading to an updated exploration graph $G^{t+1}=G^{t}\oplus(a^{t},i^{t+1})$, where $\oplus$ denotes adding a new node and edge to the graph. The Q-value function is updated accordingly, as described in Section 3.1.1.
@@
Given $a^{t}_{i}$ in $A_{\text{env}}(i^{t})$, the $Q(S^{t},a^{t}_{i})$ should be relatively low if executing it leads to a $i^{t+1}$ that is repeated in $S^{t}$, however, as mentioned in Section 3.1.2, the $i^{t+1}$ unknown at $S^{t}$. To overcome this problem, We propose a Q-value function that approximates $Q(S^{t},a^{t}_{i})$ with $Q(S^{x},a^{t}_{i})$ when there is a $S^{x}$, $x\in[1,\dots,t-1]$ with the same $a^{t}_{i}$ was executed. As for $a^{t}_{i}$ that is never executed in the past states, $Q(S^{t},a^{t}_{i})$ is set at a default value of 100.
@@
To offset the potential error in the Q-value function, we leverage the knowledge in the MLLM to determine the most suitable action. After $A^{\prime}_{\text{env}}(i^{t})$ is sampled, we let the MLLM to predict $\hat{Q}(S^{t},a^{t}_{x})$ for each sampled candidate action $a^{t}_{x}\in A^{\prime}_{\text{env}}(i^{t})$. $\hat{Q}(a^{t}_{x})$ is an MLLM-approximation of the $Q(S^{t},a^{t}_{x})$, and the in-context learning is employed. Specifically, the MLLM is provided with the current state $S^{t}$ and two example actions $(a_{\text{eg1}},a_{\text{eg2}})$ along with their corresponding Q-values $(Q(S^{t},a_{\text{eg1}}),Q(S^{t},a_{\text{eg2}}))$. The example actions are chosen from $A_{\text{env}}(i^{t})$ with $(Q(S^{t},a_{\text{eg1}}),Q(S^{t},a_{\text{eg2}}))\neq 100$ while having the corresponding element $e_{\text{eg1}}$ and $e_{\text{eg2}}$ most similar to the $e^{t}_{x}$. We develop a GUI Element Fuzzy Visual Matching module, detailed in Appendix A, to identify the visual similarity between element visuals. Next, we mark $(a_{\text{eg1}},a_{\text{eg2}})\text{ and }a^{t}_{x}$ on the screenshot of $i^{t}$ as the visual input to the MLLM with the Set-of-Mark method Yang et al. (2023). Figure 3 shows an example of this process, and the full prompts are detailed in Appendix D. If no suitable $a_{\text{eg}}$ is found, the example-related content is omitted from the prompt. Finally, the action $a^{t}_{x}$ with the highest predicted $\hat{Q}(a^{t}_{x})$ is selected as the action $a^{t}$ to be executed.
@@ 3.2 Autonomous Data Annotation with GUI-Bee
After the exploration, for each edge $a^{t}$ in the $G$, the connected nodes $(i^{t},i^{t+1})$ is sent to the MLLM to generate $u^{t}$, a list of queries serving as action grounding queries for the target element $e^{t}$ in ${i}^{t}$. We carefully design the prompt, detailed in Appendix B, to guide the MLLM in generating these queries. The queries involve both "what is currently visible" on the screen and "what will appear" after interacting with GUI elements, which we refer as System 1 and System 2 grounding queries, inspired by Kahneman (2011). Building on the multi-lens prompting method Fan et al. (2024), we create separate visual prompts, or lenses: two capture the full screens of $i^{t}$ and $i^{t+1}$ while one isolates $e^{t}$ to ensure high-quality outputs. The center of $e^{t}$ is then sampled as the target point $p^{t}$, and the resulting data ($u^{t},{i}^{t},p^{t}$) is used to fine-tune GUI action grounding models.
@@ 4 NovelScreenSpot Benchmark
NovelScreenSpot simulates real-world GUI model deployment scenarios, where App owners want to evaluate GUI action models on their own specific GUI environments, potentially novel environments that the models are not previously trained on. Therefore, unlike existing benchmarks that emphasize diversity across many environments, such as the ScreenSpot Cheng et al. (2024), NovelScreenSpot instead provides a greater data variation within each environment. Notably, the NovelScreenSpot includes a large number of queries, around one-third of the total benchmark, focusing on interaction outcomes. These queries require environment-specific knowledge and hardly exist in the existing benchmarks. The five GUI environments in the NovelScreenSpot are three offline websites from the VisualWebArena Koh et al. (2024a)—Shopping, Classifieds (a second-hand marketplace), and Reddit (an online forum)—and two online websites, Photoshop-web and Eventbrite. These environments vary greatly in style, with Photoshop-web dominated by professional icons, Shopping and Classifieds emphasizing images, and Reddit and Eventbrite focusing more on textual content.
@@ 5.1 Alignment to Novel GUI Environments
We employ the GUI-Bee agent to explore the five environments in NovelScreenSpot, conducting up to 400 exploration steps per environment with three candidate actions sampled at each step. To ensure diverse screen data, the exploration is repeated three times per environment at varying screen resolutions. In this work, we adopt GPT-4o OpenAI (2024) as the multimodal large language model (MLLM), though other MLLMs can also be integrated into the framework. The resulting exploration statistics are summarized in Table 1. Using the data generated from these explorations, we continuously fine-tune four GUI grounding models: SeeClick Cheng et al. (2024), Qwen-GUI Chen et al. (2024), and UIX-7B Liu et al. (2024) to adapt to novel environments within NovelScreenSpot. Additionally, for the alignment of Qwen-GUI and UIX-7B, we leverage both vision-only and Vision+A11y data, as these models were previously fine-tuned to accommodate similar formats. Further details on exploration configurations, data formatting, and fine-tuning settings are provided in Appendix E.
@@
We present the model performance before and after alignment to novel GUI environments in Table 2. The models are either continuously fine-tuned using exploration data collected by the GUI-Bee agent or, as a baseline, training data from Multimodal-Mind2Web Zheng et al. (2024b). Our results show that data collected by GUI-Bee significantly outperforms the baseline in improving model performance across all tested models. We attribute this to the environment-specific knowledge captured by the GUI-Bee agent, which contrasts with Multimodal-Mind2Web’s design for generalization across environments. Notably, the UIX-7B model achieves the most substantial accuracy gains and the highest overall accuracy after alignment. Furthermore, models fine-tuned in the vision+A11y GUI action grounding setting demonstrate greater improvements. This is likely because the original GUI grounding models were not trained with accessibility (A11y) tree information, and the inclusion of A11y data provides additional screen context, enabling more accurate predictions. These findings underscore the value of the diverse formats in GUI-Bee’s collected data, which incorporate contextual information such as accessibility data to enhance model performance.
@@
We extend our evaluation by following Gou et al. (2024) to leverage the test data from Multimodal-Mind2Web Zheng et al. (2024b) for assessing the alignment of GUI grounding models to novel environments. This test data simulates GUI grounding tasks guided by MLLM-planned instructions in GUI agent applications. Multimodal-Mind2Web data is derived from real action sequences across diverse GUI environments, corresponding to high-level, real-world GUI tasks. The grounding queries consist of element descriptions generated by MLLMs, without explicitly referencing coordinates. As shown in Table 2, the result reveals that models continuously fine-tuned with GUI-Bee data achieve significantly greater performance improvements compared to baseline models using Multimodal-Mind2Web training set data. This further validates the effectiveness of our proposed method.
@@
Model Performance on Action Outcomes Related Queries
@@ 5.2 Exploration Efficiency Evaluation
To further evaluate the efficiency of our GUI-Bee agent in exploring and generating data within GUI environments, we task it, along with two baseline agents, to explore three offline GUI environments: Shopping, Classifieds, and Reddit—from the Visual Web Arena Koh et al. (2024a). These environments are reset to identical initial states at the beginning of each exploration, ensuring that all agents start from the same conditions and face equivalent challenges.
@@
To evaluate the Q-ICRL method adopted by our GUI-Bee agent, we introduce two ablated version of GUI-Bee: one explore with the In-Context Reinforcement Learning (ICRL) method and the other one explore with the random strategy. The ICRL agent is an ablated version of our GUI-Bee agent. It uses the same RL state representation as the GUI-Bee agent but does not build or utilize the Q-value function. Instead, it directly leverages the MLLM with in-context learning to select the next action on the GUI screen. The in-context example is randomly chosen from the actions recorded in the exploration graph. The random agent follows a purely stochastic strategy, selecting actions randomly from the set of candidate actions on each screen. All agents are constrained to the same maximum number of action steps to enable a fair comparison.